### v1

# PEneo: Unifying Line Extraction, Line Grouping, and Entity Linking for End-to-end Document Pair Extraction

Abstract
Document pair extraction aims to identify key and value entities as well as their relationships from visually-rich documents. Most existing methods divide it into two separate tasks: semantic entity recognition (SER) and relation extraction (RE). However, simply concatenating SER and RE serially can lead to severe error propagation, and it fails to handle cases like multi-line entities in real scenarios. To address these issues, this paper introduces a novel framework, PEneo (Pair Extraction new decoder option), which performs document pair extraction in a unified pipeline, incorporating three concurrent sub-tasks: line extraction, line grouping, and entity linking. This approach alleviates the error accumulation problem and can handle the case of multi-line entities. Furthermore, to better evaluate the model’s performance and to facilitate future research on pair extraction, we introduce RFUND, a re-annotated version of the commonly used FUNSD and XFUND datasets, to make them more accurate and cover realistic situations. Experiments on various benchmarks demonstrate PEneo’s superiority over previous pipelines, boosting the performance by a large margin (e.g., 19.89%-22.91% F1 score on RFUND-EN) when combined with various backbones like LiLT and LayoutLMv3, showing its effectiveness and generality. Codes and the new annotations will be open to the public.
1 Introduction
Document pair extraction is a vital step in analyzing form-like documents containing information organized as key-value pairs. It involves identifying the key and value entities, as well as their linking relationships from document images. Previous research has generally divided it into two document understanding tasks: semantic entity recognition (SER) and relation extraction (RE). The SER task involves extracting contents that belong to predefined categories, such as retrieving store names and prices from receipts [11] or analyzing nutrition facts labels [14]. Most of the existing methods [26, 28, 27, 10, 23, 16] implement SER using BIO tagging, where tokens in the input text sequence are tagged as the beginning (B), inside (I), or outside (O) element for each entity. On the other hand, the RE task aims to identify relations between given entities, such as predicting the linkings between form elements [7, 29]. Previous works [28, 27, 23] have typically employed a linking classification network for relation extraction: given the entities in the document, it first generates representations for all possible entity pairs, then applies binary classification to filter out the valid ones. Document pair extraction is usually achieved by serially concatenating the above two tasks (SER+RE), where the SER model first identifies all the key and value entities from the document, and the RE model finds the matching values for each key. Figure 1 provides examples of SER, RE, and document pair extraction.
Although achievements have been made in SER and RE, the existing SER+RE approach overlooks several issues. In previous settings, SER and RE are viewed as two distinct tasks that have inconsistent input/output forms and employ simplified evaluation metrics. For the SER task, entity-level OCR results are usually given [7], where text lines belonging to the same entity are aggregated and serialized in human reading order. The model categorizes each token based on the well-organized sequence, neglecting the impact of improper OCR outputs. In the RE task, models take the ground truths of the SER task as input, using prior knowledge of entity content and category. The model simply needs to predict the linkings based on the provided key and value entities, and the linking-level F1 score is taken as the evaluation metric. In real-world applications, however, the situation is considerably more complex. Commonly used OCR engines typically generate results at the line level. For entities with multiple lines, an extra line grouping step is required before BIO tagging, which is hard to realize for complex layout documents. Additionally, errors in SER predictions can significantly impact the RE step, resulting in unsatisfactory pair extraction results. Section 5.5 analyzes the SER+RE performance drop in detail.
To tackle the aforementioned challenges, we propose PEneo (Pair Extraction new decoder option) to implement pair extraction in a joint manner. Our framework begins by acquiring multi-modal representations of each token using an existing document understanding backbone. Then, a newly designed decoder concurrently performs the following three sub-tasks: (1) line extraction, which identifies the text lines belonging to the key and value entities; (2) line grouping, where lines within an entity are merged; (3) entity linking, which establishes the connections between keys and their corresponding values. The three tasks are optimized jointly to minimize their discrepancies and reduce error accumulation. Subsequently, a linking parsing module integrates the output from each sub-task to generate the key-value pairs. This approach effectively suppresses errors in local predictions and produces optimal results. Notably, the decoder can collaborate with any BERT-like document understanding backbone and can be fine-tuned to downstream datasets directly without additional task-specific pre-training.
Furthermore, we found that some annotations in the two commonly used form understanding datasets, FUNSD [7] and XFUND [29], do not meet the real-world requirements well. Hence we propose their relabeled version, RFUND. The inconsistent labeling granularity in the original annotations is unified into line-level, aiming to imitate the output of a real OCR engine. We also rectified the category and relation annotations to make it more clear.
Our main contributions can be summarized as follows:
We propose PEneo, a novel framework that unifies document pair extraction through joint modeling of line extraction, line grouping, and entity linking. The model has an enhanced error suppression capability and is able to cope with challenges like multi-line entities.
We relabel the widely used FUNSD and XFUND datasets to better simulate real-world conditions for document pair extraction, including line-level OCR and more accurate annotations. The relabeled dataset is termed RFUND.
Experiments on various benchmarks show that PEneo significantly outperforms existing pipelines when collaborating with different backbones, demonstrating the effectiveness and versatility of the proposed method.
2 Related Work
Early studies [25, 22] utilized heuristic rules to extract key-value pairs from documents. These approaches exhibit limited applicability to specific document layouts and demonstrate poor generalization performance. In recent years, with the advancements in deep learning techniques, researchers have proposed several deep learning-based methods for pair extraction. LayoutLM [26] has first proposed embedding the coordinate of each word into a BERT-style model to capture the multi-modal features of each token. It also strengthens the model’s representational capability with specially designed pre-training tasks. Subsequent works primarily focus on improving the backbones to obtain more powerful and general token representations. [8] introduces a novel relative spatial encoding to capture layout information effectively. [28, 10, 16, 15, 1, 5] incorporate visual features and enhance the interaction of different modalities through newly designed architectures and pre-training tasks. [23] proposes a two-branch structure that allows flexible switching of semantic encoding modules and fast adaptation in different language scenarios. To handle the text serialization problem, ERNIE-Layout [21] employs a Layout-Parser as well as a reading order prediction task to recognize and sort the text segments. To boost performance on RE, GeoLayoutLM [19] proposes a novel relation head and a geometric pre-training schema, and obtains outstanding performance on various RE benchmarks [29, 20]. ESP [30] introduces an end-to-end pipeline incorporating text detection, text recognition, entity extraction, and entity linking. It also predicts inter/intra-linkings between words to cope with multi-line entities. The model shows outstanding performance on various relation extraction tasks. However, all of the aforementioned methods achieve pair extraction by simply concatenating the downstream SER and RE model, thereby overlooking the error accumulations in this process.
In addition to the SER+RE pipeline, some other approaches explore alternative ways to accomplish pair extraction. FUDGE [3] employs a graph-based scheme that iteratively aggregates text lines and predicts the key-value linking. Although it is capable of handling multi-line entities, its performance is relatively limited due to the absence of a powerful multi-modal encoder. SPADE [12] handles the document parsing task using the dependency parsing strategy, and it demonstrates good performance on the CORD [20] dataset. However, it models and decodes the document at the word level and thus has a huge computational overhead. Donut [13] and Dessurt [4] both propose image-to-sequence pipelines, which achieve pair extraction in a question-answering manner; QGN [2] introduces a query-driven network that generates the pair predictions using the value prefix. These generative models require a lot of training data and fail to handle complex layout documents. DocTr [17] first identifies anchor words from the input OCR results, then predicts entity-level bounding boxes and relations using a vision-language decoder, achieving structured information extraction and multi-line entity grouping. However, it requires task-specific pre-training and cannot be directly applied to existing backbones. KVPFormer [9] takes a different approach by first identifying key entities in the given documents and then employing an answer prediction module to determine their corresponding values. Notably, KVPFormer achieves exceptional performance in the RE task without pre-training, but it requires prior knowledge of entity spans and cannot handle unordered OCR inputs.
3 The RFUND Dataset
3.1 Analysis of the Original Annotations
The FUNSD dataset is a commonly used form understanding benchmark that comprises scanned English documents. The XFUND dataset is its multilingual extension, covering 7 languages (Chinese, Japanese, Spanish, French, Italian, German, and Portuguese). Entities in these forms are categorized into four types, including header, question, answer, and other. Linking relationships between different entities are annotated to represent the structure of the form.
While most contents in FUNSD are annotated at the entity level, multi-line entities with first-line indentation are annotated separately, as illustrated in Figure 1(a). XFUND exhibits variable granularity in annotations, with some contents labeled at the entity level and others at the line level, as shown in Figure 1(b). Such inconsistent labeling standards can hinder model training and fail to work in real-world scenarios. Moreover, it is observed that certain entities in both FUNSD and XFUND have category labels that differ from human understanding (Figure 1(c)), highlighting the need for refinement of the annotations in this aspect.
Lang | Split | # of entities | # of multi-line entities | # of pairs |
EN | train | 7049 | 631 | 3023 |
test | 2201 | 277 | 848 | |
ZH | train | 9948 | 1139 | 3887 |
test | 3469 | 435 | 1414 | |
JA | train | 9775 | 778 | 2875 |
test | 3390 | 342 | 1094 | |
ES | train | 11109 | 521 | 4022 |
test | 3354 | 180 | 1186 | |
FR | train | 8680 | 307 | 3444 |
test | 3499 | 153 | 1404 | |
IT | train | 11720 | 581 | 5111 |
test | 3769 | 207 | 1635 | |
DE | train | 8177 | 575 | 3500 |
test | 2645 | 202 | 1086 | |
PT | train | 11259 | 591 | 4211 |
test | 4101 | 179 | 1593 |
3.2 Relabeling for Real-World Scenarios
Given the original entity-level annotations, we employed rules to generate line-level text strings and bounding boxes. For cases that could not be handled perfectly, we made manual corrections. We then grouped text lines belonging to the same entity, including the cases of first-line indent described above. Any redundant linking labels were removed. We also rectified the entity category labels and key-value linkings to make them consistent with human understanding. We also eliminated header-to-question linkings that describe nested information, simplifying the task scope. We term the resulting dataset as RFUND, and Table 1 summarizes its key statistics.
3.3 Task Definition
The model is expected to take the line-level OCR results, which are commonly used in real scenarios, as input. It should then predict all the key-value pairs in string format. Pair-level F1 score is employed for performance comparison, where a pair prediction will be considered as True Positive if and only if both the predicted key and value text strings match the ground truths exactly.
4 PEneo
The architecture of our proposed framework is depicted in Figure 3. PEneo provides a new decoder for the pair extraction task. It first derives token representations using existing multi-modal document understanding backbones like LayoutLMv2. The decoder then concurrently generates relation matrices for three sub-tasks—line extraction, line grouping, and entity linking. Finally, a linking parsing algorithm is applied to obtain the predicted key-value pairs. We elaborate on each module in the following sections.
4.1 Multi-modal Encoder
The encoder tokenizes the input text lines into tokens and integrates semantic, layout, and visual (optional) information to obtain multi-modal features for each token. Various BERT-like document understanding models, including LayoutLMv2 [28], LayoutLMv3 [10], and LiLT [23], can serve as the encoding backbone for PEneo. To reduce memory consumption for the following operation, we append a linear projection layer to map the channels of the output features to a smaller size at the final stage:
where is the backbone output feature, is the channel reduced feature, and are parameters of the projection layer. is the output size of the backbone, and is the reduced channel size.
4.2 Joint Extraction Decoder
The decoder is responsible for both entity recognition and linking prediction. Specifically, it performs the following operations: (1) Line Extraction: Identifies the text lines belonging to the key and value entities. (2) Line Grouping: Merges lines within an entity to create cohesive representations. (3) Entity Linking: Establishes connections between key and value entities. These operations are optimized jointly to minimize discrepancies and reduce error accumulation, ensuring the overall effectiveness of PEneo.
Inspired by [24], token representations from the encoder are concatenated in a pair-wise manner. Subsequently, a pair encoding layer is applied to obtain the token pair representations matrix , where is the number of input tokens. Each entry is computed as
Here denotes vector concatenation. and are parameters of the pair encoding layer.
The matrix is then fed into three separate branches, performing line extraction, line grouping, and entity linking tasks in parallel.
4.2.1 Line Extraction
This branch extracts lines belonging to key and value entities through span prediction. A classifier is applied to to get the line extraction score :
The prediction matrix is obtained through argmax operation on , identifying the start and end tokens of lines that pertain to key or value entities, in which entries are defined as
As shown in Figure 3, the element at row Name, column Name indicates that a text line consists of a single token Name is extracted. While element at row 22, column 25 indicates that tokens 22, 56, and 25 form a target line 225625.
4.2.2 Line Grouping
To aggregate lines belonging to the same entity, we create a line head grouping matrix and a line tail grouping matrix to represent the connections between the tokens at the beginning and end of each line, respectively. For two neighboring lines within an entity, whose tokens range from and , their connections are represented by and . In Figure 3, entity 202-0-921, Sha Pei Ze Dist. consists of two text lines 202-0-921, and Sha Pei Ze Dist.. Linking predictions between their head tokens 202 and Sh, as well as their tail tokens , and . indicate that these two lines should be grouped.
4.2.3 Entity Linking
This branch predicts linkings between key and value entities. Two classifiers are applied to , forming the entity head linking matrix and the entity tail linking matrix . For a key-value pair where the key ranges from tokens and the value ranges from tokens , we have and . If the key/value entity contains multiple text lines, we establish a connection from the head token of the key entity’s first line to the head token of the value entity’s first line, as well as a connection from the tail token of the key entity’s last line to the tail token of the value entity’s last line. As shown in Figure 3, the key entity Address should be connected to value 202-0-921, Sha Pei Ze Dist.. Hence the connections between the head tokens of their first lines (Address and 202) and the tail tokens of their last lines (Address and .) are predicted as positive.
4.2.4 Linking Parsing
Matrix predicts the first token of each key and value entity. The Span of the entity’s first line can be determined by referring to the line extraction result from . For multi-line entities, based on the first and last token of each line, spans of the entity’s following lines can be retrieved iteratively from the line grouping predictions and . Once we collect all the contents of the current pair, we compare the last token of the key and value entity with the entity tail linking result to determine the validity of the current prediction. During the parsing process, if the predictions of different matrices are found to be contradictory, the current parsed content is considered to be erroneous and is directly eliminated. The algorithm flow of the linking parsing module is discussed in the supplementary materials.
4.3 Supervised Learning Target
For each prediction matrix, we adopt a weighted cross-entropy loss as its supervised learning target:
where is the prediction score matrix of each branch, is the corresponding label. is the class weighting tensor.
The overall loss of PEneo during the training phase is the weighted sum of losses from the five matrices.
where stands for the line extraction loss, and stand for the line head and tail grouping losses, and stand for the entity head and tail linking losses, are loss weighting hyper-parameters.
5 Experiments
5.1 Datasets
We conduct experiments on RFUND and SIBR [30]. As described in section 3, RFUND contains 8 subsets corresponding to 8 different languages. We follow the language-specific fine-tuning settings in [27] to evaluate the model’s performance on each subset. SIBR is a bilingual dataset composed of 600 training samples and 400 testing samples. It contains 600 Chinese invoices, 300 English bills of entry, and 100 bilingual receipts. The dataset is annotated at line level, with entity linking (inter-links) and line grouping (intra-links) labels provided. We observed some contradictory annotations in SIBR and made manual corrections. In our experiment, we focus on the linkings between question and answer entities in SIBR and employ pair-level F1 score as the evaluation metric.
5.2 Implementation Details
The reduced feature channel size is set to . During the training phase, the loss weighting hyper-parameters are all set to 1. To address category imbalance, the class weighting tensor for the cross-entropy loss is set to . We employ AdamW [18] as the optimizer. The learning rate is set to 2e-6 for the encoder backbone and 1e-4 for the decoder, scheduled by a linear scheduler with a warm-up ratio of 0.1. We fine-tune PEneo for 650 epochs on RFUND and 330 epochs on SIBR, with a batch size of 4.
5.3 Baseline Settings
We employed several widely used and publicly available models, LiLT, LayoutLMv2, LayoutXLM, and LayoutLMv3, as the encoding backbone to evaluate the effectiveness of our proposed PEneo framework. The baseline method serially combines an SER model and a RE model for pair extraction. At the SER stage, we sorted the input lines with Augmented XY Cut [6], aiming to maximize the adjacency of lines within an entity. The model learns to extract entities based on the sorted sequence and group the correctly ordered lines through BIO tagging. For the RE part, we train the model using entity-level annotations, consistent with previous studies. Additionally, we considered GeoLayoutLM [19], which includes a powerful RE decoder, as the baseline. We perform pair extraction by concatenating its downstream SER and RE models, following the aforementioned settings. Furthermore, we adopt Donut [13], an OCR-free end-to-end model with an image-to-sequence pipeline, to our pair extraction task. It takes the document image as input and predicts HTML-like strings representing the document structure. It is worth noting that some models only provide pre-trained weights for a single or a small number of languages, which does not cover all the samples in RFUND. Hence, we only evaluated the model’s performance on the language subset covered by their pre-training corpus. Further details are discussed in supplementary materials.
Method | Pipeline | F1 |
Donut [13] | End-to-End | 24.54 |
GeoLayoutLM [19] | SER+RE | 69.03 |
LiLT[EN-R] [23] | SER+RE | 54.33 |
PEneo-LiLT[EN-R] | Joint | 74.22 (+19.89) |
LiLT[InfoXLM] [23] | SER+RE | 52.18 |
PEneo-LiLT[InfoXLM] | Joint | 74.29 (+22.11) |
LayoutXLM [27] | SER+RE | 52.98 |
PEneo-LayoutXLM | Joint | 74.25 (+21.27) |
LayoutLMv2 [28] | SER+RE | 49.06 |
PEneo-LayoutLMv2 | Joint | 71.97 (+22.91) |
LayoutLMv3 [10] | SER+RE | 57.66 |
PEneo-LayoutLMv3 | Joint | 79.27 (+21.61) |
Method | Pipeline | F1 |
Donut [23] | End-to-End | 17.26 |
LiLT[InfoXLM] [23] | SER+RE | 72.76 |
PEneo-LiLT[InfoXLM] | Joint | 82.36 (+9.60) |
LayoutXLM [27] | SER+RE | 70.45 |
PEneo-LayoutXLM | Joint | 82.23 (+11.78) |
LayoutLMv3 [10] | SER+RE | 73.51 |
PEneo-LayoutLMv3 | Joint | 82.52 (+9.01) |
Method | ZH | JA | ES | FR | IT | DE | PT |
Donut [13] | 28.21 | 13.82 | - | - | - | - | - |
LayoutLMv3 [10] | 72.14 | - | - | - | - | - | - |
PEneo-LayoutLMv3 | 85.05 (+12.91) | - | - | - | - | - | - |
LiLT[InfoXLM] [23] | 66.50 | 43.98 | 63.85 | 62.60 | 60.57 | 55.13 | 52.96 |
PEneo-LiLT[InfoXLM] | 80.51 (+14.01) | 54.59 (+10.61) | 71.43 (+7.58) | 77.49 (+14.89) | 73.62 (+13.05) | 70.11 (+14.98) | 71.43 (+18.47) |
LayoutXLM [27] | 64.11 | 40.21 | 66.75 | 67.98 | 63.04 | 58.77 | 59.79 |
PEneo-LayoutXLM | 80.41 (+16.30) | 52.81 (+12.60) | 74.56 (+7.81) | 78.11 (+10.13) | 75.17 (+12.13) | 74.06 (+15.29) | 70.81 (+11.02) |
5.4 Comparison with Existing Methods
Results are shown in Table 2-4. Previous pipelines underperform on the pair extraction task. Donut utilizes the visual modality only, which may be insufficient for analyzing the diverse and complex key-value relationships in the document. We argue that Donut performs better in scenarios with a small number of output tokens, such as receipt understanding in CORD [20]. Its prediction ability is not yet fully developed for documents with a large number of texts. The SER+RE pipelines suffer from performance drop, mainly due to the error accumulation between modules and the improper text order, which will be discussed in the subsequent sections. PEneo, on the other hand, substantially improves the performance of each backbone. On RFUND-EN, the F1 score of the entire pipeline is boost by 19.89% for LiLT[EN-R], 22.11% for LiLT[InfoXLM], 21.27% for LayoutXLM, 22.91% for LayoutLMv2, and 21.61% for LayoutLMv3. Most of these backbones outperform the strong baseline GeoLayoutLM which contains task-specific pre-trained RE modules, although there exist huge gaps between them under the previous SER+RE setting. For the other language subset of RFUND, PEneo still offers substantial performance improvements. On the SIBR dataset, the new pipeline has demonstrated a score improvement ranging from 9.01% to 11.78%, confirming its ability in bilingual settings. These outcomes underscore the effectiveness and versatility of PEneo, as it consistently achieves performance gains across multiple language scenarios and diverse backbone configurations.
5.5 Analysis of Module Collaboration
Performance Drop in the SER+RE Pipeline
For the SER+RE pipeline, although the downstream models may work well on the SER or RE task, the performance drops drastically when it comes to the pair extraction setting. Figure 4 visualizes the failure cases. Erroneous predictions in the SER step greatly confuse the RE module, leading to redundant or missing output. The imperfect line ordering generated by the preprocessing step also makes it difficult to group entity lines through BIO tagging.
We observe that LiLT and LayoutLMv3 underperform in several cases, which seems to be contradictory to the results reported in previous literature [23, 10] at first glance. In fact, these two methods utilize entity-level boxes for layout modeling, while the conventional settings ([26, 28, 27, 19]) use word-level boxes. In our experiment, all the models take the line-level coordinates as input, which affects their SER ability to some extent.
To further illustrate the performance drop phenomenon in the SER+RE pipeline, we conducted experiments on RFUND-EN with LiLT[InfoXLM] using different SER results. As shown in Table 5, a performance gap of 19 points in the SER module results in a 15-point decrease in pair extraction performance (#1 and #2). As the SER score is further reduced by 3 points, the performance of the pair extraction reduces by 3 points as well (#2 and # 3, #4). Results show that the SER+RE approach exhibits a serious accumulation of errors. The accuracy of the SER module greatly affects the performance of the whole pipeline.
We then explore the impact of different types of SER errors on the whole pipeline. Commonly seen errors in the SER step include: (1) entity false negative, where keys and values are categorized as background elements; (2) entity false positive, where background elements are categorized as keys or values; (3) entity category error, where keys/values are identified as values/keys; (4) entity fragmentation, where an entity is recognized as multiple parts belong to different categories, as shown in Figure 3(a). We added the above disturbances to the SER ground truths with different probabilities and tested the performance variation of pair extraction under the setting of a fixed RE module. Results are shown in Figure 5. Entity fragmentation has the largest effect since inaccurate entity spans inherently lead to mistakes. False positives have the smallest influence as the RE module could filter out some misclassifications given accurate spans. False negatives directly cause missing links, while category errors interfere with the RE module’s reasoning. Overall, precise entity span detection proved critical for the SER+RE pipeline. However, it is highly dependent on the correct order of input text and large granularity of coordinate information [16], which is often difficult to achieve in practice, making the model underperform.
# | SER | RE | Pair Extraction |
1 | 100.00 | 67.18 | 67.18 |
2 | 80.28 (-19.72) | 67.18 | 52.18 (-15.00) |
3 | 79.20 (-20.80) | 67.18 | 51.09 (-16.09) |
4 | 76.88 (-23.12) | 67.18 | 48.42 (-18.76) |
Effectiveness of PEneo
Compared with the SER+RE scheme, our approach suppresses the error accumulation between modules, reduces the influence brought by other factors to different components, and the capacity of the backbone is fully exploited. Figure 4 shows how PEneo suppresses the errors. In Figure 3(a), although the line extraction module of PEneo initially raised some background elements, they were filtered out by referring to the line grouping and entity linking predictions. In Figure 3(b), PEneo gives correct results through the cooperation of different modules, while the SER+RE method gives false positive predictions. In Figure 3(c), the SER+RE pipeline fails to group the multi-line entity IND/LOR VOLUME in the table header, leading to erroneous predictions. PEneo, on the other hand, successfully addresses this challenge.
Encoder | Line Extraction | Line Grouping | Pair Extraction |
LiLT-I | 87.68 | 53.87 | 74.29 |
100.00 (+12.32) | 53.87 | 74.93 (+0.64) | |
87.68 | 100.00 (+46.13) | 77.14 (+2.85) | |
100.00 (+12.32) | 100.00 (+46.13) | 78.85 (+4.56) | |
LaLM3B | 92.84 | 63.44 | 79.27 |
100.00 (+7.16) | 63.44 | 80.18 (+0.91) | |
92.84 | 100.00 (+36.56) | 82.25 (+2.98) | |
100.00 (+7.16) | 100.00 (+36.56) | 83.44 (+4.17) |
To further illustrate the advantages of PEneo, we replace the predictions of line extraction and line grouping with ground truths to test for variations in pair extraction performance. As shown in Table 6, in contrast to the SER+RE pipeline, using true labels in line extraction and line grouping only led to minor pair extraction gains, despite huge gaps between the performance of the predictions and ground truths. This demonstrates PEneo’s ability to suppress downstream errors, thanks to the joint modeling and evidence accumulation of the three sub-tasks. Incorrect local predictions can be effectively rectified during the final linking parsing stage.
6 Conclusion and Future Work
In this paper, we proposed PEneo, a novel framework for end-to-end document pair extraction from visually-rich documents. By unifying the line extraction, line grouping, and entity linking tasks into a joint pipeline, PEneo effectively addressed the error propagation and challenges associated with multi-line entities. Experiments show that the proposed method outperforms previous pipelines by a large margin when collaborating with various backbones, demonstrating its effectiveness and versatility. Additionally, we introduced RFUND, a re-annotated version of the widely used FUNSD and XFUND datasets, to provide a more accurate and practical evaluation in real-world scenarios. Future work will focus on improving robustness to imperfect OCR results and complex structure parsing to enhance applicability to real-world documents. Investigating techniques like multi-task learning across the sub-tasks could further improve joint modeling. Overall, we hope this work will spark further research beyond the realms of prevalent SER+RE pipelines, and we believe the proposed PEneo provides an important step towards unified, real-world document pair extraction.

### v2

@@
 PEneo: Unifying Line Extraction, Line Grouping, and Entity Linking for End-to-end Document Pair Extraction
-Abstract
+Abstract.
 Document pair extraction aims to identify key and value entities as well as their relationships from visually-rich documents.
@@
 Experiments on various benchmarks demonstrate PEneo’s superiority over previous pipelines, boosting the performance by a large margin (e.g., 19.89%-22.91% F1 score on RFUND-EN) when combined with various backbones like LiLT and LayoutLMv3, showing its effectiveness and generality.
-Codes and the new annotations will be open to the public.
-1 Introduction
+Codes and the new annotations are available at https://github.com/ZeningLin/PEneo.
+1. Introduction
 Document pair extraction is a vital step in analyzing form-like documents containing information organized as key-value pairs.
@@
 Previous research has generally divided it into two document understanding tasks: semantic entity recognition (SER) and relation extraction (RE).
-The SER task involves extracting contents that belong to predefined categories, such as retrieving store names and prices from receipts [11] or analyzing nutrition facts labels [14].
-Most of the existing methods [26, 28, 27, 10, 23, 16] implement SER using BIO tagging, where tokens in the input text sequence are tagged as the beginning (B), inside (I), or outside (O) element for each entity.
-On the other hand, the RE task aims to identify relations between given entities, such as predicting the linkings between form elements [7, 29].
-Previous works [28, 27, 23] have typically employed a linking classification network for relation extraction: given the entities in the document, it first generates representations for all possible entity pairs, then applies binary classification to filter out the valid ones.
+The SER task involves extracting contents that belong to predefined categories, such as retrieving store names and prices from receipts (Huang et al., 2019) or analyzing nutrition facts labels (Kuang et al., 2023).
+Most of the existing methods (Xu et al., 2020, 2021b, 2021a; Huang et al., 2022; Wang et al., 2022; Li et al., 2021b) implement SER using BIO tagging, where tokens in the input text sequence are tagged as the beginning (B), inside (I), or outside (O) element for each entity.
+On the other hand, the RE task aims to identify relations between given entities, such as predicting the linkings between form elements (Guillaume Jaume, 2019; Xu et al., 2022).
+Previous works (Xu et al., 2021b, a; Wang et al., 2022) have typically employed a linking classification network for relation extraction: given the entities in the document, it first generates representations for all possible entity pairs, then applies binary classification to filter out the valid ones.
 Document pair extraction is usually achieved by serially concatenating the above two tasks (SER+RE), where the SER model first identifies all the key and value entities from the document, and the RE model finds the matching values for each key.
@@
 In previous settings, SER and RE are viewed as two distinct tasks that have inconsistent input/output forms and employ simplified evaluation metrics.
-For the SER task, entity-level OCR results are usually given [7], where text lines belonging to the same entity are aggregated and serialized in human reading order.
+For the SER task, entity-level OCR results are usually given (Guillaume Jaume, 2019), where text lines belonging to the same entity are aggregated and serialized in human reading order.
 The model categorizes each token based on the well-organized sequence, neglecting the impact of improper OCR outputs.
@@
 Notably, the decoder can collaborate with any BERT-like document understanding backbone and can be fine-tuned to downstream datasets directly without additional task-specific pre-training.
-Furthermore, we found that some annotations in the two commonly used form understanding datasets, FUNSD [7] and XFUND [29], do not meet the real-world requirements well.
+Furthermore, we found that some annotations in the two commonly used form understanding datasets, FUNSD (Guillaume Jaume, 2019) and XFUND (Xu et al., 2022), do not meet the real-world requirements well.
 Hence we propose their relabeled version, RFUND.
@@
 Experiments on various benchmarks show that PEneo significantly outperforms existing pipelines when collaborating with different backbones, demonstrating the effectiveness and versatility of the proposed method.
-2 Related Work
-Early studies [25, 22] utilized heuristic rules to extract key-value pairs from documents.
+2. Related Work
+2.1. Document Pair Extraction Methods
+Early studies (Watanabe et al., 1995; Seki et al., 2007) utilized heuristic rules to extract key-value pairs from documents.
 These approaches exhibit limited applicability to specific document layouts and demonstrate poor generalization performance.
 In recent years, with the advancements in deep learning techniques, researchers have proposed several deep learning-based methods for pair extraction.
-LayoutLM [26] has first proposed embedding the coordinate of each word into a BERT-style model to capture the multi-modal features of each token.
+LayoutLM (Xu et al., 2020) has first proposed embedding the coordinate of each word into a BERT-style model to capture the multi-modal features of each token.
 It also strengthens the model’s representational capability with specially designed pre-training tasks.
 Subsequent works primarily focus on improving the backbones to obtain more powerful and general token representations.
-[8] introduces a novel relative spatial encoding to capture layout information effectively.
-[28, 10, 16, 15, 1, 5] incorporate visual features and enhance the interaction of different modalities through newly designed architectures and pre-training tasks.
-[23] proposes a two-branch structure that allows flexible switching of semantic encoding modules and fast adaptation in different language scenarios.
-To handle the text serialization problem, ERNIE-Layout [21] employs a Layout-Parser as well as a reading order prediction task to recognize and sort the text segments.
-To boost performance on RE, GeoLayoutLM [19] proposes a novel relation head and a geometric pre-training schema, and obtains outstanding performance on various RE benchmarks [29, 20].
-ESP [30] introduces an end-to-end pipeline incorporating text detection, text recognition, entity extraction, and entity linking.
+(Hong et al., 2022) introduces a novel relative spatial encoding to capture layout information effectively.
+(Xu et al., 2021b; Huang et al., 2022; Li et al., 2021b, a; Appalaraju et al., 2021; Gu et al., 2021) incorporate visual features and enhance the interaction of different modalities through newly designed architectures and pre-training tasks.
+(Wang et al., 2022) proposes a two-branch structure that allows flexible switching of semantic encoding modules and fast adaptation in different language scenarios.
+To handle the text serialization problem, ERNIE-Layout (Peng et al., 2022) employs a Layout-Parser as well as a reading order prediction task to recognize and sort the text segments.
+To boost performance on RE, GeoLayoutLM (Luo et al., 2023) proposes a novel relation head and a geometric pre-training schema, and obtains outstanding performance on various RE benchmarks (Xu et al., 2022; Park et al., 2019).
+ESP (Yang et al., 2023) introduces an end-to-end pipeline incorporating text detection, text recognition, entity extraction, and entity linking.
 It also predicts inter/intra-linkings between words to cope with multi-line entities.
@@
 In addition to the SER+RE pipeline, some other approaches explore alternative ways to accomplish pair extraction.
-FUDGE [3] employs a graph-based scheme that iteratively aggregates text lines and predicts the key-value linking.
-Although it is capable of handling multi-line entities, its performance is relatively limited due to the absence of a powerful multi-modal encoder.
-SPADE [12] handles the document parsing task using the dependency parsing strategy, and it demonstrates good performance on the CORD [20] dataset.
-However, it models and decodes the document at the word level and thus has a huge computational overhead.
-Donut [13] and Dessurt [4] both propose image-to-sequence pipelines, which achieve pair extraction in a question-answering manner; QGN [2] introduces a query-driven network that generates the pair predictions using the value prefix.
+FUDGE (Davis et al., 2021) employs a graph-based detection scheme that iteratively aggregates text lines and predicts the key-value linking.
+Although it is capable of handling multi-line entities, its performance is relatively limited due to the absence of semantic information.
+SPADE (Hwang et al., 2021) handles the document parsing task using the dependency parsing strategy, and it demonstrates good performance on the CORD (Park et al., 2019) dataset.
+However, it models and decodes the document based on quantities of word-to-word relation, resulting in a huge computational overhead.
+Donut (Kim et al., 2022) and Dessurt (Davis et al., 2022) both propose image-to-sequence pipelines, which achieve pair extraction in a question-answering manner; QGN (Cao et al., 2022) introduces a query-driven network that generates the pair predictions using the value prefix.
 These generative models require a lot of training data and fail to handle complex layout documents.
-DocTr [17] first identifies anchor words from the input OCR results, then predicts entity-level bounding boxes and relations using a vision-language decoder, achieving structured information extraction and multi-line entity grouping.
+DocTr (Liao et al., 2023) first identifies anchor words from the input OCR results, then predicts entity-level bounding boxes and relations using a vision-language decoder, achieving structured information extraction and multi-line entity grouping.
 However, it requires task-specific pre-training and cannot be directly applied to existing backbones.
-KVPFormer [9] takes a different approach by first identifying key entities in the given documents and then employing an answer prediction module to determine their corresponding values.
-Notably, KVPFormer achieves exceptional performance in the RE task without pre-training, but it requires prior knowledge of entity spans and cannot handle unordered OCR inputs.
-3 The RFUND Dataset
-3.1 Analysis of the Original Annotations
+TPP (Zhang et al., 2023) employs a unified token path prediction framework for multi-line SER and RE tasks, and it outperforms conventional BIO-tagging baselines.
+However, TPP regards RE as a token clustering task, which leads to the inability to differentiate the key and value content individually.
+KVPFormer (Hu et al., 2023) takes a different approach by first identifying key entities in the given documents and then employing an answer prediction module to determine their corresponding values.
+Notably, its proposed spatial compatibility feature helps achieve exceptional performance in the RE task without pre-training, but it requires prior knowledge of entity spans and cannot handle unordered OCR inputs.
+2.2. Joint Extraction in Plain Texts
+Joint extraction aims to identify the subject-relation-object triplets simultaneously from plain texts.
+Mainstream approaches can be broadly categorized into two types: pipeline-based methods and joint methods.
+The pipeline-based method is akin to the SER+RE approach mentioned above.
+It involves a sequential combination of SER and RE tasks, wherein the entity contents are initially predicted, followed by the classification of the semantic relation type between them (Wang et al., 2019; Soares et al., 2019; Zhong and Chen, 2021; Ye et al., 2022).
+The joint methods unify the two tasks through specifically designed network architecture.
+(Zheng et al., 2017; Dai et al., 2019) propose novel tagging schemes that incorporate relation extraction annotations into BIO tags, thereby unifying the entire pipeline in BIO tagging manners.
+(Wei et al., 2020) employs span prediction to extract the subject content from the token sequence, subsequently predicting their corresponding objects and relation types through relation-specific object taggers.
+(Wang et al., 2020) leverages span prediction for entity identification, then constructs head and tail linking matrices for relation extraction.
+This approach enables the parsing of triplets through a joint decoding schema.
+It is noteworthy that while there are valuable insights to be gained from joint extraction in plain texts, the extraction of key-value pairs in visually-rich documents presents unique challenges that necessitate additional efforts.
+For instance, the entities to be extracted may span across multiple OCR boxes, giving rise to challenges related to multi-line SER.
+Moreover, the determination of relationships between entities relies on spatial information, thus requiring the development of specialized modules to effectively address these requirements.
+3. The RFUND Dataset
+3.1. Analysis of the Original Annotations
 The FUNSD dataset is a commonly used form understanding benchmark that comprises scanned English documents.
@@
 Entities in these forms are categorized into four types, including header, question, answer, and other.
-Linking relationships between different entities are annotated to represent the structure of the form.
-While most contents in FUNSD are annotated at the entity level, multi-line entities with first-line indentation are annotated separately, as illustrated in Figure 1(a).
+Entity-level and word-level OCR results are provided, and linking relationships between different entities are annotated to represent the structure of the form.
+While most contents in FUNSD are annotated at the entity level, multi-line entities with first-line indentation are annotated in a distinct manner.
+As illustrated in Figure 2(a), the first line of the answer paragraph is considered a separate entity, while the other lines remain aggregated as another.
+The question entity is linked to both answers, leading to redundant annotations.
 XFUND exhibits variable granularity in annotations, with some contents labeled at the entity level and others at the line level, as shown in Figure 2(b).
@@
 Moreover, it is observed that certain entities in both FUNSD and XFUND have category labels that differ from human understanding (Figure 2(c)), highlighting the need for refinement of the annotations in this aspect.
+3.2. Relabeling for Real-World Scenarios
+Based on the original entity-level and word-level OCR results, we implemented a set of rules to divide paragraphs into line-level text strings and bounding boxes.
+Specifically, we examined the vertical distance between two adjacent words within an entity.
+If the difference exceeds the average height of entity words, we assign the latter word to the subsequent line.
+For cases that could not be handled perfectly by the rules, we made manual corrections.
+To represent entity-level information, line grouping annotations were added to indicate the correct order of line aggregation.
+Cases of first-line indent described above were corrected, and any redundant linking labels were removed.
+To ensure consistency with human understanding, we adjusted the entity category labels and key-value linkings accordingly.
+In addition, we eliminated header-to-question linkings that describe nested information, simplifying the task scope.
+We term the resulting dataset as RFUND, and Table 1 summarizes its key statistics.
+3.3. Task Definition
+The model is expected to take the line-level OCR results, which are commonly used in real scenarios, as input.
+It should then predict all the key-value pairs in string format.
+Pair-level F1 score is employed for performance comparison, where a pair prediction will be considered as True Positive if and only if both the predicted key and value text strings exactly match the ground truths.
 Lang | Split | # of entities | # of multi-line entities | # of pairs |
@@
 test | 4101 | 179 | 1593 |
-3.2 Relabeling for Real-World Scenarios
-Given the original entity-level annotations, we employed rules to generate line-level text strings and bounding boxes.
-For cases that could not be handled perfectly, we made manual corrections.
-We then grouped text lines belonging to the same entity, including the cases of first-line indent described above.
-Any redundant linking labels were removed.
-We also rectified the entity category labels and key-value linkings to make them consistent with human understanding.
-We also eliminated header-to-question linkings that describe nested information, simplifying the task scope.
-We term the resulting dataset as RFUND, and Table 1 summarizes its key statistics.
-3.3 Task Definition
-The model is expected to take the line-level OCR results, which are commonly used in real scenarios, as input.
-It should then predict all the key-value pairs in string format.
-Pair-level F1 score is employed for performance comparison, where a pair prediction will be considered as True Positive if and only if both the predicted key and value text strings match the ground truths exactly.
-4 PEneo
+4. PEneo
 The architecture of our proposed framework is depicted in Figure 3.
@@
 We elaborate on each module in the following sections.
-4.1 Multi-modal Encoder
+4.1. Multi-modal Encoder
 The encoder tokenizes the input text lines into tokens and integrates semantic, layout, and visual (optional) information to obtain multi-modal features for each token.
-Various BERT-like document understanding models, including LayoutLMv2 [28], LayoutLMv3 [10], and LiLT [23], can serve as the encoding backbone for PEneo.
+Various BERT-like document understanding models, including LayoutLMv2 (Xu et al., 2021b), LayoutLMv3 (Huang et al., 2022), and LiLT (Wang et al., 2022), can serve as the encoding backbone for PEneo.
 To reduce memory consumption for the following operation, we append a linear projection layer to map the channels of the output features to a smaller size at the final stage:
 where is the backbone output feature, is the channel reduced feature, and are parameters of the projection layer. is the output size of the backbone, and is the reduced channel size.
-4.2 Joint Extraction Decoder
+4.2. Joint Extraction Decoder
 The decoder is responsible for both entity recognition and linking prediction.
@@
 These operations are optimized jointly to minimize discrepancies and reduce error accumulation, ensuring the overall effectiveness of PEneo.
-Inspired by [24], token representations from the encoder are concatenated in a pair-wise manner.
+Inspired by (Wang et al., 2020), token representations from the encoder are concatenated in a pair-wise manner.
 Subsequently, a pair encoding layer is applied to obtain the token pair representations matrix , where is the number of input tokens.
@@
 The matrix is then fed into three separate branches, performing line extraction, line grouping, and entity linking tasks in parallel.
-4.2.1 Line Extraction
+4.2.1. Line Extraction
 This branch extracts lines belonging to key and value entities through span prediction. A classifier is applied to to get the line extraction score :
@@
 While element at row 22, column 25 indicates that tokens 22, 56, and 25 form a target line 225625.
-4.2.2 Line Grouping
+4.2.2. Line Grouping
 To aggregate lines belonging to the same entity, we create a line head grouping matrix and a line tail grouping matrix to represent the connections between the tokens at the beginning and end of each line, respectively.
@@
 indicate that these two lines should be grouped.
-4.2.3 Entity Linking
+4.2.3. Entity Linking
 This branch predicts linkings between key and value entities.
@@
 Hence the connections between the head tokens of their first lines (Address and 202) and the tail tokens of their last lines (Address and .) are predicted as positive.
-4.2.4 Linking Parsing
+4.2.4. Linking Parsing
 Matrix predicts the first token of each key and value entity.
@@
 During the parsing process, if the predictions of different matrices are found to be contradictory, the current parsed content is considered to be erroneous and is directly eliminated.
-The algorithm flow of the linking parsing module is discussed in the supplementary materials.
-4.3 Supervised Learning Target
+4.3. Supervised Learning Target
 For each prediction matrix, we adopt a weighted cross-entropy loss as its supervised learning target:
@@
 where stands for the line extraction loss, and stand for the line head and tail grouping losses, and stand for the entity head and tail linking losses, are loss weighting hyper-parameters.
-5 Experiments
-5.1 Datasets
-We conduct experiments on RFUND and SIBR [30].
+5. Experiments
+5.1. Datasets
+We conduct experiments on RFUND and SIBR (Yang et al., 2023).
 As described in section 3, RFUND contains 8 subsets corresponding to 8 different languages.
-We follow the language-specific fine-tuning settings in [27] to evaluate the model’s performance on each subset.
+We follow the language-specific fine-tuning settings in (Xu et al., 2021a) to evaluate the model’s performance on each subset.
 SIBR is a bilingual dataset composed of 600 training samples and 400 testing samples.
@@
 In our experiment, we focus on the linkings between question and answer entities in SIBR and employ pair-level F1 score as the evaluation metric.
-5.2 Implementation Details
+5.2. Implementation Details
 The reduced feature channel size is set to .
@@
 To address category imbalance, the class weighting tensor for the cross-entropy loss is set to .
-We employ AdamW [18] as the optimizer.
+We employ AdamW (Loshchilov and Hutter, 2019) as the optimizer.
 The learning rate is set to 2e-6 for the encoder backbone and 1e-4 for the decoder, scheduled by a linear scheduler with a warm-up ratio of 0.1.
 We fine-tune PEneo for 650 epochs on RFUND and 330 epochs on SIBR, with a batch size of 4.
-5.3 Baseline Settings
+5.3. Baseline Settings
 We employed several widely used and publicly available models, LiLT, LayoutLMv2, LayoutXLM, and LayoutLMv3, as the encoding backbone to evaluate the effectiveness of our proposed PEneo framework.
 The baseline method serially combines an SER model and a RE model for pair extraction.
-At the SER stage, we sorted the input lines with Augmented XY Cut [6], aiming to maximize the adjacency of lines within an entity.
-The model learns to extract entities based on the sorted sequence and group the correctly ordered lines through BIO tagging.
+At the SER stage, we sorted the input lines with Augmented XY Cut (Gu et al., 2022), aiming to maximize the adjacency of lines within an entity.
+The model learns to extract entities based on the sorted sequence and group the correctly ordered lines through BIO tagging(Ramshaw and Marcus, 1999).
 For the RE part, we train the model using entity-level annotations, consistent with previous studies.
-Additionally, we considered GeoLayoutLM [19], which includes a powerful RE decoder, as the baseline.
-We perform pair extraction by concatenating its downstream SER and RE models, following the aforementioned settings.
-Furthermore, we adopt Donut [13], an OCR-free end-to-end model with an image-to-sequence pipeline, to our pair extraction task.
-It takes the document image as input and predicts HTML-like strings representing the document structure.
+Additionally, we adapted FUDGE (Davis et al., 2021), Donut (Kim et al., 2022), GeoLayoutLM (Luo et al., 2023), TPP (Zhang et al., 2023) and GPT-4V (Achiam et al., 2023) to the pair extraction task.
+FUDGE employs a graph-based detection pipeline that predicts the key-value bounding box pairs.
+Since it only predicts boxes, we report its performance on the box-pair F1-score as a compromise.
+Donut takes the document image as input and predicts HTML-like strings containing key-value pairs.
+GeoLayoutLM is a strong baseline that includes a powerful RE decoder, and we perform pair extraction by concatenating its downstream SER and RE models, following the aforementioned SER+RE settings.
+TPP employs a token clustering scheme, and the groups of pair tokens can be obtained by parsing its VrD-EL matrix with depth-first searching.
+Its performance is reported on token-group F1-score.
+For GPT-4V, we follow the evaluation pipeline proposed by (Shi et al., 2023).
 It is worth noting that some models only provide pre-trained weights for a single or a small number of languages, which does not cover all the samples in RFUND.
 Hence, we only evaluated the model’s performance on the language subset covered by their pre-training corpus.
-Further details are discussed in supplementary materials.
-Method | Pipeline | F1 |
-Donut [13] | End-to-End | 24.54 |
-GeoLayoutLM [19] | SER+RE | 69.03 |
-LiLT[EN-R] [23] | SER+RE | 54.33 |
-PEneo-LiLT[EN-R] | Joint | 74.22 (+19.89) |
-LiLT[InfoXLM] [23] | SER+RE | 52.18 |
-PEneo-LiLT[InfoXLM] | Joint | 74.29 (+22.11) |
-LayoutXLM [27] | SER+RE | 52.98 |
-PEneo-LayoutXLM | Joint | 74.25 (+21.27) |
-LayoutLMv2 [28] | SER+RE | 49.06 |
-PEneo-LayoutLMv2 | Joint | 71.97 (+22.91) |
-LayoutLMv3 [10] | SER+RE | 57.66 |
-PEneo-LayoutLMv3 | Joint | 79.27 (+21.61) |
-Method | Pipeline | F1 |
-Donut [23] | End-to-End | 17.26 |
-LiLT[InfoXLM] [23] | SER+RE | 72.76 |
-PEneo-LiLT[InfoXLM] | Joint | 82.36 (+9.60) |
-LayoutXLM [27] | SER+RE | 70.45 |
-PEneo-LayoutXLM | Joint | 82.23 (+11.78) |
-LayoutLMv3 [10] | SER+RE | 73.51 |
-PEneo-LayoutLMv3 | Joint | 82.52 (+9.01) |
+Method | Venue | Pipeline | F1 |
+FUDGE (Davis et al., 2021) | ICDAR’21 | End-to-End | 53.15‡ |
+Donut (Kim et al., 2022) | ECCV’22 | Image2Seq | 24.54 |
+GeoLayoutLM (Luo et al., 2023) | CVPR’23 | SER+RE | 69.03 |
+TPP-LayoutLMv3BASE (Zhang et al., 2023) | EMNLP’23 | Joint | 50.27‡ |
+GPT-4V w/o OCR (Achiam et al., 2023) | arXiv’23 | Image2Seq | 20.96 |
+GPT-4V w OCR (Achiam et al., 2023) | arXiv’23 | Image2Seq | 38.15 |
+LiLT[EN-R]BASE (Wang et al., 2022) | ACL’22 | SER+RE | 54.33 |
+PEneo-LiLT[EN-R]BASE | Ours | Joint | 74.22 (+19.89) |
+LiLT[InfoXLM]BASE (Wang et al., 2022) | ACL’22 | SER+RE | 52.18 |
+PEneo-LiLT[InfoXLM]BASE | Ours | Joint | 74.29 (+22.11) |
+LayoutXLMBASE (Xu et al., 2021a) | arXiv’21 | SER+RE | 52.98 |
+PEneo-LayoutXLMBASE | Ours | Joint | 74.25 (+21.27) |
+LayoutLMv2BASE (Xu et al., 2021b) | ACL’21 | SER+RE | 49.06 |
+PEneo-LayoutLMv2BASE | Ours | Joint | 71.97 (+22.91) |
+LayoutLMv3BASE (Huang et al., 2022) | ACMMM’22 | SER+RE | 57.66† |
+PEneo-LayoutLMv3BASE | Ours | Joint | 79.27 (+21.61) |
+Method | Venue | Pipeline | F1 |
+DonutBASE (Wang et al., 2022) | ECCV’22 | Image2Seq | 17.26 |
+LiLT[InfoXLM]BASE (Wang et al., 2022) | ACL’22 | SER+RE | 72.76 |
+PEneo-LiLT[InfoXLM]BASE | Ours | Joint | 82.36 (+9.60) |
+LayoutXLMBASE (Xu et al., 2021a) | arXiv’21 | SER+RE | 70.45 |
+PEneo-LayoutXLMBASE | Ours | Joint | 82.23 (+11.78) |
+LayoutLMv3 (Huang et al., 2022) | ACMMM’22 | SER+RE | 73.51† |
+PEneo-LayoutLMv3 | Ours | Joint | 82.52 (+9.01) |
 Method | ZH | JA | ES | FR | IT | DE | PT |
-Donut [13] | 28.21 | 13.82 | - | - | - | - | - |
-LayoutLMv3 [10] | 72.14 | - | - | - | - | - | - |
+DonutBASE (Kim et al., 2022) | 28.21 | 13.82 | - | - | - | - | - |
+LayoutLMv3 (Huang et al., 2022) | 72.14† | - | - | - | - | - | - |
 PEneo-LayoutLMv3 | 85.05 (+12.91) | - | - | - | - | - | - |
-LiLT[InfoXLM] [23] | 66.50 | 43.98 | 63.85 | 62.60 | 60.57 | 55.13 | 52.96 |
-PEneo-LiLT[InfoXLM] | 80.51 (+14.01) | 54.59 (+10.61) | 71.43 (+7.58) | 77.49 (+14.89) | 73.62 (+13.05) | 70.11 (+14.98) | 71.43 (+18.47) |
-LayoutXLM [27] | 64.11 | 40.21 | 66.75 | 67.98 | 63.04 | 58.77 | 59.79 |
-PEneo-LayoutXLM | 80.41 (+16.30) | 52.81 (+12.60) | 74.56 (+7.81) | 78.11 (+10.13) | 75.17 (+12.13) | 74.06 (+15.29) | 70.81 (+11.02) |
-5.4 Comparison with Existing Methods
+LiLT[InfoXLM]BASE (Wang et al., 2022) | 66.50 | 43.98 | 63.85 | 62.60 | 60.57 | 55.13 | 52.96 |
+PEneo-LiLT[InfoXLM]BASE | 80.51 (+14.01) | 54.59 (+10.61) | 71.43 (+7.58) | 77.49 (+14.89) | 73.62 (+13.05) | 70.11 (+14.98) | 71.43 (+18.47) |
+LayoutXLMBASE (Xu et al., 2021a) | 64.11 | 40.21 | 66.75 | 67.98 | 63.04 | 58.77 | 59.79 |
+PEneo-LayoutXLMBASE | 80.41 (+16.30) | 52.81 (+12.60) | 74.56 (+7.81) | 78.11 (+10.13) | 75.17 (+12.13) | 74.06 (+15.29) | 70.81 (+11.02) |
+5.4. Comparison with Existing Methods
 Results are shown in Table 2-4.
 Previous pipelines underperform on the pair extraction task.
-Donut utilizes the visual modality only, which may be insufficient for analyzing the diverse and complex key-value relationships in the document.
-We argue that Donut performs better in scenarios with a small number of output tokens, such as receipt understanding in CORD [20].
+FUDGE and Donut utilize the visual modality only, which may be insufficient for analyzing the diverse and complex key-value relationships in the document.
+We argue that Donut performs better in scenarios with a small number of output tokens, such as receipt understanding in CORD (Park et al., 2019).
 Its prediction ability is not yet fully developed for documents with a large number of texts.
+TPP employs a unified token-path-prediction pipeline to tackle the RE task, without a dedicated design for error suppression.
+In this schema, even a minor error in the linking prediction matrix could result in a completely incorrect outcome.
+For GPT-4V, integrating the OCR results into the prompt can significantly improve its performance, but it still falls short of numerous supervised approaches.
+Upon analyzing its output, we argue that its underperformance can be attributed primarily to the LLM hallucination, which leads to numerous redundant or inaccurate predictions.
 The SER+RE pipelines suffer from performance drop, mainly due to the error accumulation between modules and the improper text order, which will be discussed in the subsequent sections.
 PEneo, on the other hand, substantially improves the performance of each backbone.
-On RFUND-EN, the F1 score of the entire pipeline is boost by 19.89% for LiLT[EN-R], 22.11% for LiLT[InfoXLM], 21.27% for LayoutXLM, 22.91% for LayoutLMv2, and 21.61% for LayoutLMv3.
+On RFUND-EN, the F1 score of the entire pipeline is boost by 19.89% for LiLT[EN-R]BASE, 22.11% for LiLT[InfoXLM]BASE, 21.27% for LayoutXLMBASE, 22.91% for LayoutLMv2BASE, and 21.61% for LayoutLMv3BASE.
 Most of these backbones outperform the strong baseline GeoLayoutLM which contains task-specific pre-trained RE modules, although there exist huge gaps between them under the previous SER+RE setting.
@@
 These outcomes underscore the effectiveness and versatility of PEneo, as it consistently achieves performance gains across multiple language scenarios and diverse backbone configurations.
-5.5 Analysis of Module Collaboration
+5.5. Analysis of Module Collaboration
 Performance Drop in the SER+RE Pipeline
@@
 The imperfect line ordering generated by the preprocessing step also makes it difficult to group entity lines through BIO tagging.
-We observe that LiLT and LayoutLMv3 underperform in several cases, which seems to be contradictory to the results reported in previous literature [23, 10] at first glance.
-In fact, these two methods utilize entity-level boxes for layout modeling, while the conventional settings ([26, 28, 27, 19]) use word-level boxes.
+We observe that LiLT and LayoutLMv3 underperform in several cases, which seems to be contradictory to the results reported in previous literature (Wang et al., 2022; Huang et al., 2022) at first glance.
+In fact, these two methods utilize entity-level boxes for layout modeling, while the conventional settings (Xu et al., 2020, 2021b, 2021a; Luo et al., 2023) use word-level boxes.
 In our experiment, all the models take the line-level coordinates as input, which affects their SER ability to some extent.
-To further illustrate the performance drop phenomenon in the SER+RE pipeline, we conducted experiments on RFUND-EN with LiLT[InfoXLM] using different SER results.
+To further illustrate the performance drop phenomenon in the SER+RE pipeline, we conducted experiments on RFUND-EN with LiLT[InfoXLM]BASE using different SER results.
 As shown in Table 5, a performance gap of 19 points in the SER module results in a 15-point decrease in pair extraction performance (#1 and #2).
@@
 Overall, precise entity span detection proved critical for the SER+RE pipeline.
-However, it is highly dependent on the correct order of input text and large granularity of coordinate information [16], which is often difficult to achieve in practice, making the model underperform.
+However, it is highly dependent on the correct order of input text and large granularity of coordinate information (Li et al., 2021b), which is often difficult to achieve in practice, making the model underperform.
 # | SER | RE | Pair Extraction |
@@
 Incorrect local predictions can be effectively rectified during the final linking parsing stage.
-6 Conclusion and Future Work
+6. Conclusion and Future Work
 In this paper, we proposed PEneo, a novel framework for end-to-end document pair extraction from visually-rich documents.
@@
 Overall, we hope this work will spark further research beyond the realms of prevalent SER+RE pipelines, and we believe the proposed PEneo provides an important step towards unified, real-world document pair extraction.
+Acknowledgements.
+This research is supported in part by National Natural Science Foundation of China (Grant No.: 62441604, 61936003).References
+- (1)
+- Achiam et al.
+(2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, et al.
+2023.
+Gpt-4 technical report.
+arXiv preprint arXiv:2303.08774 (2023).
+- Appalaraju et al.
+(2021) Srikar Appalaraju, Bhavan Jasani, Bhargava Urala Kota, Yusheng Xie, and R Manmatha.
+2021.
+DocFormer: End-to-End Transformer for Document Understanding.
+In 2021 IEEE/CVF International Conference on Computer Vision (ICCV).
+IEEE, 973–983.
+- Cao et al.
+(2022) Haoyu Cao, Xin Li, Jiefeng Ma, Deqiang Jiang, Antai Guo, Yiqing Hu, Hao Liu, Yinsong Liu, and Bo Ren.
+2022.
+Query-Driven Generative Network for Document Information Extraction in the Wild.
+In Proceedings of the 30th ACM International Conference on Multimedia.
+4261–4271.
+- Dai et al.
+(2019) Dai Dai, Xinyan Xiao, Yajuan Lyu, Shan Dou, Qiaoqiao She, and Haifeng Wang.
+2019.
+Joint Extraction of Entities and Overlapping Relations Using Position-Attentive Sequence Labeling.
+In Proceedings of the AAAI Conference on Artificial Intelligence, Vol.
+33.
+6300–6308.
+- Davis et al.
+(2022) Brian Davis, Bryan Morse, Brian Price, Chris Tensmeyer, Curtis Wigington, and Vlad Morariu.
+2022.
+End-to-End Document Recognition and Understanding with Dessurt.
+In European Conference on Computer Vision.
+Springer, 280–296.
+- Davis et al.
+(2021) Brian Davis, Bryan Morse, Brian Price, Chris Tensmeyer, and Curtis Wiginton.
+2021.
+Visual FUDGE: Form Understanding via Dynamic Graph Editing.
+In Document Analysis and Recognition–ICDAR 2021: 16th International Conference, Lausanne, Switzerland, September 5–10, 2021, Proceedings, Part I 16.
+Springer, 416–431.
+- Gu et al.
+(2021) Jiuxiang Gu, Jason Kuen, Vlad I Morariu, Handong Zhao, Rajiv Jain, Nikolaos Barmpalios, Ani Nenkova, and Tong Sun.
+2021.
+UniDoc: Unified Pretraining Framework for Document Understanding.
+Advances in Neural Information Processing Systems 34 (2021), 39–50.
+- Gu et al.
+(2022) Zhangxuan Gu, Changhua Meng, Ke Wang, Jun Lan, Weiqiang Wang, Ming Gu, and Liqing Zhang.
+2022.
+XYLayoutLM: Towards Layout-aware Multimodal Networks for Visually-rich Document Understanding.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+4583–4592.
+- Guillaume Jaume (2019) Jean-Philippe Thiran Guillaume Jaume, Hazim Kemal Ekenel. 2019. FUNSD: A Dataset for Form Understanding in Noisy Scanned Documents. In ICDAR-OST.
+- Hong et al.
+(2022) Teakgyu Hong, Donghyun Kim, Mingi Ji, Wonseok Hwang, Daehyun Nam, and Sungrae Park.
+2022.
+BROS: A Pre-trained Language Model Focusing on Text and Layout for Better Key Information Extraction from Documents.
+In Proceedings of the AAAI Conference on Artificial Intelligence, Vol.
+36.
+10767–10775.
+- Hu et al.
+(2023) Kai Hu, Zhuoyuan Wu, Zhuoyao Zhong, Weihong Lin, Lei Sun, and Qiang Huo.
+2023.
+A Question-Answering Approach to Key Value Pair Extraction from Form-Like Document Images.
+Proceedings of the AAAI Conference on Artificial Intelligence 37, 11 (Jun.
+2023), 12899–12906.
+- Huang et al.
+(2022) Yupan Huang, Tengchao Lv, Lei Cui, Yutong Lu, and Furu Wei.
+2022.
+LayoutLMv3: Pre-training for Document AI with Unified Text and Image Masking.
+In Proceedings of the 30th ACM International Conference on Multimedia.
+4083–4091.
+- Huang et al.
+(2019) Zheng Huang, Kai Chen, Jianhua He, Xiang Bai, Dimosthenis Karatzas, Shijian Lu, and CV Jawahar.
+2019.
+ICDAR2019 Competition on Scanned Receipt OCR and Information Extraction.
+In 2019 International Conference on Document Analysis and Recognition (ICDAR).
+IEEE, 1516–1520.
+- Hwang et al.
+(2021) Wonseok Hwang, Jinyeong Yim, Seunghyun Park, Sohee Yang, and Minjoon Seo.
+2021.
+Spatial Dependency Parsing for Semi-Structured Document Information Extraction.
+In Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021.
+330–343.
+- Kim et al.
+(2022) Geewook Kim, Teakgyu Hong, Moonbin Yim, JeongYeon Nam, Jinyoung Park, Jinyeong Yim, Wonseok Hwang, Sangdoo Yun, Dongyoon Han, and Seunghyun Park.
+2022.
+OCR-free Document Understanding Transformer.
+In European Conference on Computer Vision.
+Springer, 498–517.
+- Kuang et al.
+(2023) Jianfeng Kuang, Wei Hua, Dingkang Liang, Mingkun Yang, Deqiang Jiang, Bo Ren, and Xiang Bai.
+2023.
+Visual information extraction in the wild: practical dataset and end-to-end solution.
+In International Conference on Document Analysis and Recognition.
+Springer, 36–53.
+- Li et al.
+(2021a) Peizhao Li, Jiuxiang Gu, Jason Kuen, Vlad I Morariu, Handong Zhao, Rajiv Jain, Varun Manjunatha, and Hongfu Liu.
+2021a.
+SelfDoc: Self-Supervised Document Representation Learning.
+In 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR).
+IEEE, 5648–5656.
+- Li et al.
+(2021b) Yulin Li, Yuxi Qian, Yuechen Yu, Xiameng Qin, Chengquan Zhang, Yan Liu, Kun Yao, Junyu Han, Jingtuo Liu, and Errui Ding.
+2021b.
+StrucTexT: Structured Text Understanding with Multi-modal Transformers.
+In Proceedings of the 29th ACM International Conference on Multimedia.
+1912–1920.
+- Liao et al.
+(2023) Haofu Liao, Aruni RoyChowdhury, Weijian Li, Ankan Bansal, Yuting Zhang, Zhuowen Tu, Ravi Kumar Satzoda, R Manmatha, and Vijay Mahadevan.
+2023.
+DocTr: Document transformer for structured information extraction in documents.
+In Proceedings of the IEEE/CVF International Conference on Computer Vision.
+19584–19594.
+- Loshchilov and Hutter (2019) Ilya Loshchilov and Frank Hutter. 2019. Decoupled Weight Decay Regularization. In Proceedings of the 7th International Conference on Learning Representations (ICLR).
+- Luo et al.
+(2023) Chuwei Luo, Changxu Cheng, Qi Zheng, and Cong Yao.
+2023.
+GeoLayoutLM: Geometric Pre-training for Visual Information Extraction.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+7092–7101.
+- Park et al.
+(2019) Seunghyun Park, Seung Shin, Bado Lee, Junyeop Lee, Jaeheung Surh, Minjoon Seo, and Hwalsuk Lee.
+2019.
+CORD: A Consolidated Receipt Dataset for Post-OCR Parsing.
+In Workshop on Document Intelligence at NeurIPS 2019.
+- Peng et al.
+(2022) Qiming Peng, Yinxu Pan, Wenjin Wang, Bin Luo, Zhenyu Zhang, Zhengjie Huang, Yuhui Cao, Weichong Yin, Yongfeng Chen, Yin Zhang, et al.
+2022.
+ERNIE-Layout: Layout Knowledge Enhanced Pre-training for Visually-rich Document Understanding.
+In Findings of the Association for Computational Linguistics: EMNLP 2022.
+3744–3756.
+- Ramshaw and Marcus (1999) Lance A Ramshaw and Mitchell P Marcus. 1999. Text chunking using transformation-based learning. In Natural language processing using very large corpora. Springer, 157–176.
+- Seki et al.
+(2007) Minenobu Seki, Masakazu Fujio, Takeshi Nagasaki, Hiroshi Shinjo, and Katsumi Marukawa.
+2007.
+Information Management System Using Structure Analysis of Paper/Electronic Documents and Its Applications.
+In Ninth International Conference on Document Analysis and Recognition (ICDAR 2007), Vol.
+2.
+689–693.
+https://doi.org/10.1109/ICDAR.2007.4377003
+- Shi et al.
+(2023) Yongxin Shi, Dezhi Peng, Wenhui Liao, Zening Lin, Xinhong Chen, Chongyu Liu, Yuyi Zhang, and Lianwen Jin.
+2023.
+Exploring ocr capabilities of gpt-4v (ision): A quantitative and in-depth evaluation.
+arXiv preprint arXiv:2310.16809 (2023).
+- Soares et al.
+(2019) Livio Baldini Soares, Nicholas Fitzgerald, Jeffrey Ling, and Tom Kwiatkowski.
+2019.
+Matching the Blanks: Distributional Similarity for Relation Learning.
+In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics.
+2895–2905.
+- Wang et al.
+(2019) Haoyu Wang, Ming Tan, Mo Yu, Shiyu Chang, Dakuo Wang, Kun Xu, Xiaoxiao Guo, and Saloni Potdar.
+2019.
+Extracting Multiple-Relations in One-Pass with Pre-Trained Transformers.
+In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics.
+1371–1377.
+- Wang et al.
+(2022) Jiapeng Wang, Lianwen Jin, and Kai Ding.
+2022.
+LiLT: A Simple yet Effective Language-Independent Layout Transformer for Structured Document Understanding.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers).
+7747–7757.
+- Wang et al.
+(2020) Yucheng Wang, Bowen Yu, Yueyang Zhang, Tingwen Liu, Hongsong Zhu, and Limin Sun.
+2020.
+TPLinker: Single-stage Joint Extraction of Entities and Relations Through Token Pair Linking.
+In Proceedings of the 28th International Conference on Computational Linguistics.
+1572–1582.
+- Watanabe et al.
+(1995) T.
+Watanabe, Qin Luo, and N.
+Sugie.
+1995.
+Layout Recognition of Multi-Kinds of Table-Form Documents.
+IEEE Transactions on Pattern Analysis and Machine Intelligence 17, 4 (1995), 432–445.
+https://doi.org/10.1109/34.385976
+- Wei et al.
+(2020) Zhepei Wei, Jianlin Su, Yue Wang, Yuan Tian, and Yi Chang.
+2020.
+A Novel Cascade Binary Tagging Framework for Relational Triple Extraction.
+In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics.
+1476–1488.
+- Xu et al.
+(2020) Yiheng Xu, Minghao Li, Lei Cui, Shaohan Huang, Furu Wei, and Ming Zhou.
+2020.
+LayoutLM: Pre-training of Text and Layout for Document Image Understanding.
+In Proceedings of the 26th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining.
+1192–1200.
+- Xu et al.
+(2021a) Yiheng Xu, Tengchao Lv, Lei Cui, Guoxin Wang, Yijuan Lu, Dinei Florencio, Cha Zhang, and Furu Wei.
+2021a.
+LayoutXLM: Multimodal Pre-training for Multilingual Visually-rich Document Understanding.
+arXiv:2104.08836 [cs.CL]
+- Xu et al.
+(2022) Yiheng Xu, Tengchao Lv, Lei Cui, Guoxin Wang, Yijuan Lu, Dinei Florencio, Cha Zhang, and Furu Wei.
+2022.
+XFUND: A Benchmark Dataset for Multilingual Visually Rich Form Understanding.
+In Findings of the Association for Computational Linguistics: ACL 2022.
+3214–3224.
+- Xu et al.
+(2021b) Yang Xu, Yiheng Xu, Tengchao Lv, Lei Cui, Furu Wei, Guoxin Wang, Yijuan Lu, Dinei Florencio, Cha Zhang, Wanxiang Che, et al.
+2021b.
+LayoutLMv2: Multi-modal Pre-training for Visually-rich Document Understanding.
+In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers).
+2579–2591.
+- Yang et al.
+(2023) Zhibo Yang, Rujiao Long, Pengfei Wang, Sibo Song, Humen Zhong, Wenqing Cheng, Xiang Bai, and Cong Yao.
+2023.
+Modeling Entities as Semantic Points for Visual Information Extraction in the Wild.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+15358–15367.
+- Ye et al.
+(2022) Deming Ye, Yankai Lin, Peng Li, and Maosong Sun.
+2022.
+Packed Levitated Marker for Entity and Relation Extraction.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers).
+4904–4917.
+- Zhang et al.
+(2023) Chong Zhang, Ya Guo, Yi Tu, Huan Chen, Jinyang Tang, Huijia Zhu, Qi Zhang, and Tao Gui.
+2023.
+Reading Order Matters: Information Extraction from Visually-rich Documents by Token Path Prediction.
+In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing.
+13716–13730.
+- Zheng et al.
+(2017) Suncong Zheng, Feng Wang, Hongyun Bao, Yuexing Hao, Peng Zhou, and Bo Xu.
+2017.
+Joint Extraction of Entities and Relations Based on a Novel Tagging Scheme.
+In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers).
+Association for Computational Linguistics, Vancouver, Canada, 1227–1236.
+https://doi.org/10.18653/v1/P17-1113
+- Zhong and Chen (2021) Zexuan Zhong and Danqi Chen.
+2021.
+A Frustratingly Easy Approach for Entity and Relation Extraction.
+In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies.
+50–61.